OpenStudy (anonymous):

Write each equation in slope-intercept from and then find the slope and y-intercept. x-4y=8

OpenStudy (anonymous):

-4y = 8 - x 4y= x- 8 y = 1/4x - 2 so here slope is 1/4 for y-int, juz plug x = 0 and find the points

OpenStudy (anonymous):

x-8=4y y = 1/4x -2 the gradient/slope is therefore 1/4 and y-intercept is -2 (0,-2).

OpenStudy (anonymous):

Thanks.